In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the larger type, a printed page contains 1,800 words. Using smaller type, a page contains 2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be in smaller type?
Answer by Earlsdon(6294)

You can let L = # of pages of large type and S = # of pages of small type.
The total number of pages is: L + S = 21 Rewrite as: L = 21 - S
The total number of words is: L(1,800) + S(2,400) = 48,000
Substitute the L = 21 - S into the second equation and solve for S.
(21 - S)(1,800) + S(2,400) = 48,000 Simplify and solve for S.
37,800 - 1,800S + 2,400S = 48,000
37,800 + 600S = 48,000 Subtract 37,800 from both sides.
600S = 10,200 Divide both sides by 600.
S = 17 The number of pages of small type.
L = 21-S
L = 21-17
L = 4 The number of pages of large type.
You can check this solution by verifying the following:
4 pages (1,800 words/page) + 17 pages(2,400 words/page = 48,000 words